### v1

# The Peripatetic Hater: Predicting Movement Among Hate Subreddits

Abstract
Many online hate groups exist to disparage others based on race, gender identity, sex, or other characteristics. The accessibility of these communities allows users to join multiple types of hate groups (e.g., a racist community and misogynistic community), which calls into question whether these peripatetic users could be further radicalized compared to users that stay in one type of hate group. However, little is known about the dynamics of joining multiple types of hate groups, nor the effect of these groups on peripatetic users. In this paper, we develop a new method to classify hate subreddits, and the identities they disparage, which we use to better understand how users become peripatetic (join different types of hate subreddits). The hate classification technique utilizes human-validated LLMs to extract the protected identities attacked, if any, across 168 subreddits. We then cluster identity-attacking subreddits to discover three broad categories of hate: racist, anti-LGBTQ, and misogynistic. We show that becoming active in a user’s first hate subreddit can cause them to become active in additional hate subreddits of a different category. We also find that users who join additional hate subreddits, especially of a different category, become more active in hate subreddits as a whole and develop a wider hate group lexicon. We are therefore motivated to train an AI model that we find usefully predicts the hate categories users will become active in based on post text read and written. The accuracy of this model may be partly driven by peripatetic users often using the language of hate subreddits they eventually join. Overall, these results highlight the unique risks associated with hate communities on a social media platform, as discussion of alternative targets of hate may lead users to target more protected identities.
Introduction
Recently, there has been extensive work documenting online hate communities and the impact that they have on the users of social media platforms (Schmitz et al. 2024; Russo et al. 2023). Such research is motivated by the combination of substantial increases in hate crimes (of Investigation 2021), and the observation that radicalized individuals attribute their beliefs to online sources (Jensen et al. 2018). Much of the existing literature in this domain provides either case studies of individual communities (Chandrasekharan et al. 2017; Schmitz et al. 2024; Russo et al. 2023) or broad analyses of groups centered around a common theme, such as the manosphere (Ribeiro et al. 2021). Additionally, although prior work observes strong user overlap between anti-feminism and alt-right communities (Mamié, Horta Ribeiro, and West 2021), such research does not show how the core themes discussed in these communities are influenced by the users active in both communities.
Here, we extend the existing research corpus by curating a comprehensive dataset of 120 hate subreddits and using LLMs and clustering techniques to group them by ideology (racist, misogynistic, and anti-LGBTQ). In these subreddits, we analyze peripatetic users, who participate in multiple categories of hate subreddits, attempting to characterize them and understand the impact that they have on Reddit’s ecosystem of hate. Our research questions are as follows:
  1. RQ1
    Are users who participate in one hate community likely to participate in hate communities of other types?
  2. RQ2
    What characteristics distinguish peripatetic and non-peripatetic users?
  3. RQ3
    How might joining additional hate communities affect user behavior?
  4. RQ4
    Given that a user has already participated in a hate community, can we predict the types of hate communities in which they will participate in the future?
Our findings help answer each respective research question: (RQ1) Using a matched-pair causal model, we find that participation in a given hate subreddit is associated with joining other hate subreddits at a rate roughly seven times higher than that of control users. (RQ2) Using BERTopic (Grootendorst 2022) and lexicons of ingroup language, we find that peripatetic users introduce into their initial subreddit topics of discussion from the categories in which they eventually participate. For example, a user who initially posts only in a misogynistic subreddit, but subsequently also posts in a racist subreddit, is more likely to introduce racist topics into the misogynistic subreddit. (RQ3) We find users that join additional hate subreddits become more active in hate subreddits as a whole and use lexicons associated with these newly joined subreddits. Both results are more pronounced when users join hate subreddits of a different category then the first subreddit they joined. (RQ4) Finally, we develop a deep-learning model to predict the subreddit categories these peripatetic users will join. This task is challenging, as we are differentiating among users all of whom already exhibit signs of hostility. We find that the language of users who eventually become peripatetic is strongly associated with the subreddit categories that they subsequently join.
Overall, our results offer a novel perspective on how hate communities may converge and become more “general” in an online environment as a result of peripatetic user activity. All data and code are in the following repository https://github.com/dan-hickey1/peripatetic-hater.
Figure 1: UMAP plot of hate speech distributions for each subreddit. Each point represents a subreddit, and their color corresponds with the K-means cluster they have been assigned to.
Related Work
Types of Online Hate Communities
Misogynist. Extensive work has investigated the various identities targeted by online hate communities. One of the most widely studied groups is the manosphere, a collection of communities centered around misogyny and anti-feminism (Marwick and Caplan 2018; Habib, Srinivasan, and Nithyanand 2022). Despite sharing this common theme, there is a wide diversity of ideologies within the manosphere. Incels, for example, believe they are incapable of finding a sexual or romantic partner and blame women and society for such misfortune (Daly and Reed 2022). This is in contrast with pick-up artists, who discuss strategies for seducing women, and “men going their own way” (MGTOW), who voluntarily decide to be celibate as they believe they should distance themselves from women (Lin 2017). A recent analysis shows MGTOW and incel communities are growing in popularity and are more toxic than other manosphere communities, which are decreasing in popularity (Ribeiro et al. 2021). Critically, members of incel communities have engaged in mass shooting events (Barcellona 2022).
Anti-LGBTQ. Anti-LGBTQ communities are also prevalent on the internet, including trans-exclusionary radical feminists (TERFs), an anti-transgender group. TERFs generally subscribe to a binary view of gender that is aligned with biological sex (disregarding intersex individuals), and see transgender people as a threat to women’s spaces (Jones and Slater 2020). Scholars have previously demonstrated flaws in such discourse and the harm it produces (Williams 2020; Hotine 2021). TERF communities have been present on Reddit in the past, with r/GenderCritical being the most prominent (Tiffany 2020). Previous work has found these communities to be more toxic than mainstream feminist communities (Balci et al. 2023), and models have been built to automatically identify the discourse employed by them (Lu and Jurgens 2022).
Racist. Racist groups have been highly prevalent on the internet, and on Reddit in particular (Chandrasekharan et al. 2017), as have a uniquely online group known as the alt-right, an extreme political movement that gained popularity through its utilization of the unique characteristics of social media (Winter 2019). The alt-right believe that the cultural identity of white individuals is at risk, with movements for social justice and political correctness being seen as principal threats (Center 2023). Alt-right online communities have radicalized users in part through the use of memes (Dafaure 2020), with one analysis showing that communities linked to the alt-right have a strong influence on the spread of memes in the larger web ecosystem (Zannettou et al. 2018). The misogynist movement “GamerGate” helped the alt-right movement gain adherents (Bezio 2018), demonstrating a relationship between seemingly distinct types of hate.
Radicalization Pathways and Gateway Communities
Radicalization is largely explicable in terms of three non-exclusive effects: (1) uncertainty reduction (Hogg, Kruglanski, and van den Bos 2013), (2) significance quest (Kruglanski et al. 2014), and (3) the need to be popular (Siegel et al. 2011). There is also substantial research into the micro-scale processes that drive people to radicalization. From a community-level perspective, scholars have studied whether certain communities are “gateways” to more extreme communities, such as the connection between the intellectual dark web and the alt-right (Ribeiro et al. 2020), or between the manosphere and the alt-right (Ribeiro et al. 2020; Barcellona 2022). Gateway communities to political and conspiracy subreddits have been automatically identified using a network analysis approach (Rollo et al. 2022). Additional research on conspiracy communities demonstrates how users who increase their engagement in specific conspiracy theory communities over time also increase their participation in less specialized conspiracy theory communities (Phadke, Samory, and Mitra 2022). Other studies measure traits indicative of user radicalization or extremism that are adopted by members of certain hate groups, finding that such traits increase once users participate in hate subreddits (Habib, Srinivasan, and Nithyanand 2022; Schmitz et al. 2024). While recommendation algorithms are often blamed for radicalizing users, recent evidence shows that users who consume extreme content on YouTube are driven to do so by external sources, and they already have high racial and gender resentment before engaging in such content (Chen et al. 2023).
Movement Among Online Communities
There is a growing body of literature devoted to understanding how users participate in multiple communities on the same platform. One theme in this research is the attempt to understand inter-community conflict, where members active in one community make posts in other communities to attack them (Kumar et al. 2018; Efstratiou et al. 2023; Datta and Adar 2019). Another paper speculates that mass migrations of users among communities on Reddit may disrupt the linguistic evolution of destination communities, which could influence their likelihood of being banned (Habib et al. 2021). More generally, Rollo et al. (Rollo et al. 2022) devise a model of “attention flow” to quantitatively describe how users on Reddit shift their focus among subreddits over time. Scholars have also built models to predict whether users will participate in particular online communities (Jin, Lee, and Cheung 2010; Joyce and Kraut 2006; Debaere, Coussement, and De Ruyck 2018), with one study finding that the social support is predictive of users remaining in an online health community (Wang et al. 2017).
The present study builds on previous research by examining a multitude of hate communities while assessing the connections among all of them. In addition, we describe the tendency of users to join specific additional hate communities after initially joining a given hate community and measure the user behavior that predates such additional participation. Finally, we build a model to predict which new hate subreddits users will join.
Methods
Collecting a Dataset of Hate Subreddits
To properly understand how different categories of hate subreddits influence each other, we must first define a set of hate subreddits to analyze that spans a variety of targeted identities. While researchers have previously curated exhaustive lists of subreddits that represent a particular ideology, such as the manosphere (Ribeiro et al. 2021) or the alt-right (Mamié, Horta Ribeiro, and West 2021), there currently exists no comprehensive dataset of hate subreddits that spans the diversity of ideologies that could be observed on the platform. For this reason, we create a large dataset of hate subreddits labeled by the identities they target. Our process consists of collecting a wide set of subreddits that are likely to be hate subreddits, detecting hate speech of different types in said subreddits, and clustering subreddits together that contain hate speech of similar types.
To identify potential hate subreddits, we collect curated lists from various sources. First, we find lists of hate subreddits or banned subreddits previously identified in the literature (Mamié, Horta Ribeiro, and West 2021; Ribeiro et al. 2021; Vidgen et al. 2021a; Schmitz et al. 2024). As these papers turn to subreddits such as r/AgainstHateSubreddits to find potential hate groups, we supplement these lists with more subreddits by searching “list of hate subreddits” on r/AgainstHateSubreddits and finding each post listing hate subreddits within the first 100 search results as of October 2023. We then collect the entire history of comments and submissions from each subreddit using the Pushshift API (Baumgartner et al. 2020). All subreddits with fewer than 1000 users are removed from further analysis, as such subreddits represent less than 2% of all users in the dataset. Additionally, as many banned subreddits could be unrelated to hate (such as black market subreddits), one annotator (who is also an author of this paper) viewed 10 random comments and 10 random submissions from each subreddit, filtering out subreddits that were clearly unrelated to hate. At the end of this process, 168 subreddits remained.
After collecting an initial set of potential hate subreddits, we define various types of hate speech within each subreddit based on the identities targeted by the speech, informed by “bias categories” in United States federal hate crime law  [^1] https://www.justice.gov/hatecrimes/learn-about-hate-crimes, where a hate crime is defined as “a crime motivated by bias against race, color, religion, national origin, sexual orientation, gender, gender identity, or disability.” We deviate from these bias categories by separating the “religion” category into two categories that represent speech directed at Jews and Muslims, as hate speech directed at these groups often additionally targets their race or national origin. We also add a category representing fat-shaming, as this type of hate speech is shown to have been prominent on Reddit (Chandrasekharan et al. 2017). Therefore, our final categories of hate speech are antisemitism, Islamophobia, ableism, fat-shaming, misogyny, xenophobia, racism, homophobia, and transphobia.
To find the prevalence of each type of hate speech in each subreddit, we utilize GPT-4, as it has been shown to perform well on the task of hate speech detection (Kumarage, Bhattacharjee, and Garland 2024). For each type of hate speech, we prompt GPT-4 to count the number of posts containing that type of hate speech in batches of 10 with the following string:
Below is a collection of 10 Reddit posts that may or may not be attacking someone because of their [BIAS CATEGORY]. Please respond with the number of these posts that are very clearly [TYPE OF HATE SPEECH]. Do not include anything else in your response:
[LIST OF POSTS]
For each subreddit, we sample a total of 20 comments and 20 submissions to be submitted to GPT-4 for hate speech detection. The sample size was chosen as a trade-off between sufficiently sampling the subreddit, and avoiding excessive API costs as well as remaining within the input token limit of 128K tokens. Subreddits in which fewer than 10% of posts contain hate speech are removed from further analysis. This leaves us with a final set of 120 subreddits.
To validate our method of detecting hate speech, three annotators who are undergraduate or recently graduated students with a background in social science, annotated 10 comments and 10 submissions each from a set of 25 subreddits with which they were ethnographically familiar, having previously conducted research on the given subreddits. The annotators assessed the number of comments and submissions containing hate speech in each category for each subreddit, achieving a Krippendorff’s alpha score of 0.5 (Hayes and Krippendorff 2007), implying relatively poor inter-rater reliability. This low alpha score demonstrates the difficulty annotators have when assessing identities targeted, alike to the common issue of low annotator agreement in hatespeech classification (Vidgen et al. 2021b; Sachdeva et al. 2022). Comparing the average values of the human annotations to the estimates obtained by GPT-4, we achieve an overall $R^{2}$ value of 0.45 for our hate speech detection method. These results suggest GPT-4 is not perfect, but can still be a good proxy of human annotators.
To separate these subreddits into distinct categories, we employ $k$-means clustering. Each subreddit is assigned a vector representing the number of posts of each type of hate speech detected in the subreddit, normalized by its total number of posts containing hate speech. We run the $k$-means algorithm for values of $k$ ranging from two to 20 and choose the number of clusters that yield the highest silhouette score (three clusters). The average distributions of types of hate speech for each cluster are shown in Figure 2 and a UMAP plot of the distributions of hate for each subreddit is shown in Figure 1.
We release the set of subreddits, their clusters, and the average number of posts containing hate speech of each category for each batch of posts, available at the following URL: https://github.com/dan-hickey1/peripatetic-hater. Upon publication, we intend to adhere to FAIR principles (FORCE11 2020) by making the data publicly available and easily accessible via GitHub. The data will be in a widely used format (CSV). Using the subreddit names, data from each subreddit can be obtained through Pushshift (Baumgartner et al. 2020). Caution should be exercised when using the dataset, as several subreddit names contain offensive phrases directed toward marginalized groups.
Figure 2: Distributions of the types of hate predicted by GPT-4 in each cluster.
We label as peripatetic users those individuals who post in an alternative category of hate subreddit at some point after their initial post in a given hate subreddit. Because we seek to explore predictive markers of such peripatetic behavior after users have already participated in an antisocial environment, we restrict our focus to users who post in an alternative category of hate subreddit within six weeks of posting in their first hate subreddit, as we believe early peripatetic behavior will be more closely linked to the text expressed in the initial stages of hate subreddit participation. As the choice of this threshold is somewhat arbitrary, we compare individuals who become peripatetic users to users who do not subsequently post in additional categories of hate subreddits for all downstream analyses. For all analyses, we perform robustness checks varying this threshold between three weeks and one year.
As automated accounts are likely to be ubiquitous across categories of subreddits but are not relevant to the questions at hand, we filter out probable bots from our dataset. Our method follows Schmitz et al. 2024 and is performed by removing accounts with the following keywords: “bot,” “auto,” “transcriber,” “gif,” “link,” and “twitter.” In total, our dataset contains approximately 3M users, 117K of whom post in an alternate category of hate subreddit within six weeks of their first post in a hate subreddit, and 2.5M of whom only post in one category of hate subreddit.
User Matching
To understand whether users who initially post in hate subreddits are at risk of becoming engaged in hate subreddits of other types, we employ a process to match each user from a given hate subreddit (a “joiner”) with a similar “counterpart” user who did not engage in a hate subreddit before the joiner, as described in Schmitz et al. 2024. For a given hate subreddit, we only consider users who have posted in the subreddit as their first hate subreddit. For each hate subreddit, the process consists of finding subreddits that users from the hate subreddit are most active in relative to their total size, and then sampling potential counterparts from those subreddits. Each user from the hate subreddit is matched to a counterpart using Mahalanobis distance matching. The features used to measure the Mahalanobis distance between users are each user’s total karma, their total number of submissions, their total number of comments, their account creation date, and the total number of posts they made in the top 50 subreddits in which the members of the given hate subreddit were most active. For each counterpart user, all features were measured prior to the month the treatment user posted in the hate subreddit. From these sets of users, we calculate the mean number of subreddits of a different type than the given hate subreddit that are joined within our chosen time threshold of six weeks following the users’ first post in the given subreddit. We report the average values for each category of subreddit as a whole. As counterpart users by definition do not participate in the hateful subreddit of their treatment counterpart, the number of subreddits joined by each counterpart user is measured within the same time period as the hate subreddit user to whom they are matched.
In addition to quantifying the level of movement across subreddits through sets of matched users, for each hate subreddit category, we measure the proportion of users who originate in that category that eventually also post in a hate subreddit belonging to another category of hate.
Evaluating peripatetic users’ use of language associated with destination subreddits in their original subreddit
To explore how ingroup language spreads among subreddits, we use the Sparse Additive GEnerative (SAGE) model of text to extract keywords characteristic of each community (Eisenstein, Ahmed, and Xing 2011). To find the words most characteristic of a set of subreddits, SAGE compares the text from the subreddit to a baseline corpus. For each category of hate subreddit, we obtain the top 300 most characteristic words from SAGE, using the text from the other categories combined as a baseline corpus.
Using the lexicon of ingroup language associated with each category of hate subreddit, we seek to understand how users who post in multiple categories of hate subreddits introduce such language to different types of communities. For each user, we identify the category of hate subreddit in which they first posted. Then, in each category, we identify users who eventually posted in a different category of hate subreddit. We then compare the language employed by such users to the language used by those who exclusively posted within a single category. For peripatetic users, we only considered posts prior to posting in other hate subreddits, and all posts were collected within each user’s first three days of activity in any hate subreddit. For each combination of origin and destination subreddits, we count the number of terms used from the lexicon of the destination category relative to the number of terms used from all other categories. These are counted separately for peripatetic users and non-peripatetic users. By counting the number of terms associated with each category, we are able to calculate the odds that a peripatetic user employs terms from their destination category within the original category before they posted in the destination category, relative to the odds that non-peripatetic users would use the same terms (i.e., the odds ratio). We assess the statistical significance of changes in language use using Fisher’s exact test.
In addition to comparing the ingroup language used by peripatetic users vs. non-peripatetic users, we also measure the relative increase in ingroup language use from an alternate category after peripatetic users become active in said alternate category, as well as the percentage change in their activity rate (number of posts per week) in all hate subreddits. For each peripatetic user, a symmetric window of time is used to measure behavior before and after becoming peripatetic (e.g., if a user posts in a misogynistic subreddit three weeks after posting in a racist subreddit, their behavior in that three weeks is measured as the “before” period and their behavior three weeks after posting in the misogynistic subreddit is measured as the “after” period).
Extracting Topics of Discussion Used by Peripatetic Users
In addition to understanding how content associated with hate subreddits in which peripatetic users eventually post appears in their postings in their initial community, we are also interested in the general themes expressed by peripatetic users, unbounded by a specific source. To understand these themes, we employ topic modeling using BERTopic (Grootendorst 2022). BERTopic uses pre-trained transformer models to embed text, then clusters those embeddings to form topics. The topics are then summarized using the highest TF-IDF words for each cluster. For each hate subreddit, we fit a topic model using the first three days of posts for each user. For peripatetic users who posted in another subreddit category within the first three days, only posts prior to this event are considered. As a preprocessing step, we remove URLs and the strings “[deleted]” and “[removed]” from all posts. We use the “outlier reduction” feature of BERTopic to assign posts that were not assigned to a topic when initially fitting the model to the closest topic based on the cosine similarity of the text embeddings. We then merge the topic models of all subreddits of the same category. After merging the topic models, we count the number of posts in the top 100 topics that peripatetic users and non-peripatetic users employ. For each topic, the odds of a peripatetic user making a post that fits within the given topic are calculated relative to the odds of a non-peripatetic user making a post on the same topic. When comparing the odds ratios of topics, only topics that appear in at least ten percent of the subreddits within the given category are considered, as some topics may come from individual subreddits that are more or less likely to contain peripatetic users. The five topics with the highest and lowest log-odds ratios for each category are reported.
Predicting Participation In Hate Subreddit Types
To probe the dynamics wherein users join additional categories of hate subreddits, we initially seek to establish whether certain factors are predictive of a user joining a different category of hate subreddit in the first place. Specifically, as we establish how language spreads among the categories of hate subreddits, we are interested in how the text to which users are exposed, and the text that they themselves employ, is indicative of their participation in each type of hate subreddit. To examine the text to which users are exposed, we utilize the posts to which the user replies, as we are confident that the user has read these posts. However, it is important to note that this approach a) excludes a great deal of text that the users likely read within the given subreddit, and b) may reflect self-selection, as the text to which users reply is likely to be especially related to their interests. We frame this as a multi-label classification problem, where we build a neural network trained on the embeddings of each type of text, and have the neural network predict whether a user subsequently joins a subreddit in each category. The text each user writes or replies to in any subreddit within three days of posting in their initial hate subreddit (including their first post in that subreddit) is embedded and fed into our model as input features (we explain embedding below). To prevent data leakage, if, within the three-day window, a user posts in a hate subreddit outside of the category in which they first posted, only posts before that point are considered. As we are primarily interested in users posting in other categories of subreddits, and not continuing to post in their original category, we do not consider a user making multiple posts in their original category as a positive example. Using this strategy, we compile a dataset of approximately 100k randomly sampled users, 70% of which was used for training, with 15% being used each for validation and testing.
For each user in the dataset, all of the text that they post in their initial subreddit category is chronologically concatenated into the same document, with each post separated by the ‘$</s>$’ token. The text to which the users reply is separately concatenated in the same way. Each document is embedded into a 768-dimensional vector using Longformer (Beltagy, Peters, and Cohan 2020). The vectors are then concatenated, resulting in a 1,538-dimensional input layer for the neural network. As some data are missing from Pushshift (Gaffney and Matias 2018), there are some cases in which all of the posts to which a user replied could not be retrieved; in those cases, a single token (‘UNK’) was used as a placeholder for the parent text. The category of subreddit in which each user initially posted, as well as the categories the users to whom they replied posted prior to their interaction, are input as a one-hot encoding vector into the neural network at the second layer. The final neural network is three layers, with 774 dimensions in the second layer (where six of the input features derive from the one-hot encoding describing the initial subreddit and the subreddits of the parent users) and 384 dimensions in the third layer (i.e., each intermediate layer divides the number of features from the text vectors by a factor of two, with additional features in the second layer due to the one-hot encoding). At each hidden layer, a dropout of 0.6 is used, as well as the leaky ReLU activation function. The softmax activation function is used at the final output layer. We perform the training on an internal computing cluster using an NVIDIA RTX A6000 GPU.
In addition to training a model utilizing both the context and author vectors, we train a model utilizing each vector independently. The same activation functions and dropout values are used, and the number of dimensions is divided by the same factor within each intermediate layer.
Finally, as there is variation in the performance of the model among individual runs, we take random splits of training, validation, and testing and train the model 50 times to calculate the mean and standard errors of the performance metrics. In addition, we only report performance for each category considering users who did not originate in that category. In contrast, the performance would be trivially almost perfect (ROC-AUC$>0.95$) if we also predict whether a user will remain in their present category, as that occurs for the vast majority of users.
Results
Subreddit Clustering.
Figure 2 shows the average number of posts predicted as hate speech for each cluster identified by the $k$-means algorithm. From the distributions, it is clear that one cluster exhibits a high degree of xenophobia and racism, another cluster exhibits high homophobia and transphobia, and the last cluster is high in misogyny. Therefore, we refer to the categories of subreddits as “racist,” “anti-LGBTQ,” and “misogynistic,” respectively. We find 69 racist subreddits, 32 anti-LGBTQ subreddits, and 19 misogynistic subreddits. Figure 1 shows the distributions of average hate scores for each subreddit plotted in two-dimensional space. Observing the plot, the categories form distinct clusters. Our results are robust to different clustering methods.
RQ1: Users who participate in one type of hate subreddit are likely to subsequently participate in hate subreddits of another type.
Figure 3: Proportion of users who post in alternate subreddits within six weeks days after becoming active in their initial subreddit. Subreddits of the same category as the user’s initial category are not counted. Error bars represent standard errors.
Figure 3 displays the proportion of newcomers who post in hate subreddits with matched treatment/counterpart pairs. It is apparent that, on average, the users who post in each hate subreddit subsequently participate in many more subreddits of another type than do users who do not post in the given hate subreddit (6-8 times higher depending on the type of subreddit; unweighted mean across all three categories: 6.9). The proportion of users who originate in racist subreddits is lower than the number joined by those who originate in the other categories, likely because there are more subreddits in the racist category. Using the Wilcoxon signed-rank test to assess differences between treatment and counterpart groups for each category, we find that the observed results are extremely unlikely to happen by chance, with $p<10^{-5}$ for each result.
RQ1: The probability of posting in an alternate category of hate subreddit does not depend on the original category.
Figure 4: Heatmaps representing peripatetic user behavior (A) Proportion of users from the original subreddit type that subsequently posted in the alternate subreddit. (B) Odds ratios of peripatetic vs. non-peripatetic users using language from the alternate hate subreddit lexicon within the origin subreddit type. (C) Change in use of language from alternate hate subreddit lexicons after posting in alternate hate subreddits. (D) Percentage changes in activity rates after posting in alternate hate subreddits. Diagonals in A, C, and D represent users who do not post in alternate subreddits but post in multiple subreddits of the same category. Stars represent statistical significance ($p<0.001$)
Figure 4A displays the rate at which members who originate in each subreddit category subsequently post in other categories of subreddits. From the figure, it is immediately apparent that, for each subreddit type, the probability of ending up there as an alternate category is the same regardless of the original category. The average proportions in the columns of the heatmap (excluding cells in which the alternate and original categories are the same) are $5.5\%,0.8\%$, and $2.2\%$, which, when normalized, are almost identical to the distribution of total users that post in each category ($65.7\%$, $9.3\%$, and $25.8\%$).
RQ2: Peripatetic users introduce language from alternate hate subreddits to the hate subreddit category in which they initially posted.
Figure 4B displays the odds of peripatetic users posting material, in the category to which they first contributed, that employs language common in the peripatetic user’s subsequent subreddit category, relative to the odds that non-peripatetic users employ such language in the first category. The heatmap reveals that every result is statistically significant and is associated with peripatetic users employing more ingroup language from the destination subreddit relative to non-peripatetic users. The most prominent differences occur when the destination category is racist and the original category is misogynistic, and when the destination is misogynistic and the origin is anti-LGBTQ. Conversely, peripatetic users who eventually join misogynistic subreddits after being active in a racist subreddit exhibit a less substantial increase in misogynistic ingroup language within racist subreddits, though the result is still statistically significant.
RQ2: Peripatetic users discuss specific topics at different rates than non-peripatetic users.
Figure 5: Most frequent topics by category of subreddit. The number of posts within each topic is displayed on the x-axis of each figure. Stopwords were removed from topic representations.
Figure 5 displays the ten most common topics within each category of subreddit. Among these topics, it is apparent that types of hate other than those that form the basis for the categorization of the subreddits are discussed. For example, topics represented by the word “race,” “Muslim,” or antisemitic slurs appear in the racist category, and topics represented by the word “gay” appear in the anti-LGBTQ category. In the misogynistic category, discussions about the gender pay gap and sexual violence are among the most frequent.
Figure 6: Topics with the five highest/lowest log-odds ratios of use between peripatetic users and non-peripatetic users. Positive values indicate topics that are over-represented among peripatetic users. Topics with an adjusted p-value greater than 0.01 according to Fisher’s exact test are lighter to show statistical insignificance. Stopwords were removed from topic representations.
The topics that peripatetic users are most and least likely to address relative to non-peripatetic users are depicted for each category in Figure 6. In both racist and misogynistic subreddits, peripatetic users are more involved in discourse about “social justice warriors” (discourse that is associated with both racism and misogyny (Brock 2015)). Additionally, in misogynistic subreddits, topics represented by the term “Nazi” and “trans” are more likely to be used by peripatetic users. From qualitative inspection, the topic represented by the word “sex” in the anti-LGBTQ category seems to capture discourse from TERF subreddits. This topic is much less likely to be used by peripatetic users. This could be because the other subreddit categories are likely to express views in opposition to all branches of feminism, including the type of feminism with which TERFs identify.
RQ3: Joining additional hate communities is associated with users become more active in hate groups and developing a broader hate-group lexicon. Complementing these results, Figure 4C shows the odds ratio of ingroup language after users join other hate subreddits (either of the same or a different category), and Figure 4D shows the percent activity increase. These results show that peripatetic users have a broader hate group lexicon, and become more active in hate groups overall. We caution that these results are not causal, as we do not compare against matched users who do not join additional hate subreddits. These results nonetheless motivate a root concern: not only is joining one hate subreddit harmful (Schmitz, Burghardt, and Muric 2022; Schmitz et al. 2024), but joining additional subreddits presents a compounding harm to users and social media.
RQ4: The text that users post, and to which they reply, is predictive of whether they will post in additional categories of hate subreddits.
ROC-AUC
Category All Target Context
Racist .60 $\pm$ .002 .59 $\pm$ .002 .62 $\pm$ .002
Anti-LGBTQ .62 $\pm$ .003 .59 $\pm$ .004 .58 $\pm$ .004
Misogynistic .62 $\pm$ .003 .61 $\pm$ .003 .59 $\pm$ .003
Table 1: Model performance.
Table 1 lists the performance for our model predicting whether a user will post in a given category of subreddit. “Target” refers to posts made by the user for whom the prediction is being made, while “context” refers to posts to which they reply. The performance of each category excludes predictions for users whose initial posts were in that category. Uncertainties represent standard errors. For anti-LGBTQ and misogynistic subreddits, the model performs best when supplied with text written by the users and the text to which they replied, compared to either input type in isolation. However, the context alone is more predictive for users who originate in racist subreddits than both the context and the target.
Discussion
This study provides a comprehensive view of the user dynamics among multiple types of hate communities on Reddit. These results alone are remarkable as they suggest that hate manifested in the current historical moment among English-speaking users of this platform can be objectively categorized into three dimensions. But why three? The dimension we term “Racist” is associated with racism, xenophobia, and antisemitism, which are types of hate that stem from an essentialist point of view, or the false notion that there are inherent biological differences among people of different races or national origins (Roth, van Stee, and Regla-Vargas 2023). The second dimension, in contrast, is associated with homophobia, transphobia, and, to a lesser degree, misogyny, i.e., hate associated with perceived choice (the false belief being that being LGBTQ is not an immutable characteristic of the individual, but rather is under their control). Finally, the third category is misogyny and, to a lesser extent, fat shaming (as we observe fat shaming is often directed at women (Bound Alberti 2021), and is associated with the objectification of women), which is hate associated with gender. These categories are intrinsically disparate types of hate even though the identities attacked are diverse within each category. How this generalizes to other social media, and how it relates to psychological perceptions of hate, constitute rich avenues for future research.
We have also shown that users who initially participate in one hate community are prone to participate in additional categories of hate communities and that participating in additional categories of hate is associated with greater activity in hate communities and a broader hate group lexicon. We can, however, better predict which users will join additional communities compared to baselines. This may be in part because peripatetic users employ more of the language from within the community that they will eventually participate in. Below we discuss the possible implications of these results concerning how hate communities can influence each other online.
RQ1: users likely to participate in multiple communities of hate. Figure 3 shows that users who originate in one type of hate community are more likely to participate in an alternate type of hate community than are similar users who did not participate in the original hate community type. Moreover, the likelihood of joining a given alternate type of hate community does not differ depending on the original type (Figure 4A). Taken together, these results imply that each type of hate community serves as a “gateway” to other hate communities. In other words, the presence of each type of community could further amplify the growth of the others, as their user bases appeal to each other. Future work can trace the reasons why users from each category of hate are more likely to join other hate subreddits – for example, the impact of recruitment on peripatetic behavior can be studied (Russo, Ribeiro, and West 2023).
RQ2: characteristics distinguishing peripatetic and non-peripatetic users. Given the results displayed in Figure 4B, it is tempting to conclude that joining additional hate groups leads peripatetic users to import hate terms and ideas when posting in the subreddit in which they first became active. We explored this possibility, but could not find conclusive evidence of the causal direction. Two non-mutually-exclusive explanations therefore seem plausible. First, it is possible that users who become peripatetic were interested from the outset in the topics of the additional subreddits that they eventually joined, i.e., differences in language use between users who do and do not subsequently become peripatetic reflect differences in the breadth of their antipathies at time zero. Second, users who eventually post in other subreddits may have been consuming content from those subreddits long before they first contributed to them, hence said content may have influenced their posts in the first subreddit which they joined. Regardless of which of these various possibilities applies, given that a) social influence likely occurs, and b) consuming content from, and contributing content to, multiple communities is facilitated by their presence on a single platform, the harm posed by hosting multiple types of hate communities on one platform may be greater than the sum of the harm posed by each community. In this scenario, hate subreddits serve as recruitment venues for each other, both enhancing the growth of their respective user bases and leading to a convergence of hate beliefs over time.
RQ3: Joining additional hate communities may further radicalize users. Figure 4C–D demonstrates the change in user behavior as they join additional hate subreddits. Namely, we see an increase in activity and a broader lexicon of these users, especially if they join subreddits of a different category. This may imply that users become more active in, and internalize language of, these various dimensions of hate. These features could be a first sign that users may radicalize more quickly by joining additional hate groups (Schmitz et al. 2024).
RQ4: Predicting peripatetic movement. Table 1 illustrates our model’s prediction performance for the types of hate subreddits users that will eventually join. The relatively poor performance of the model suggests that stochastic or unobservable factors strongly influence who becomes a peripatetic hater. Nonetheless, for each category, performance is above the baseline, indicating that peripateticism is not a purely spurious decision. Users are either partly motivated to join based on their past multi-hate category opinions represented in writing, or their writing indicates that they are more vulnerable to being influenced to join other types of hate. This technology is the first step towards understanding users’ motivations for joining multiple hate subreddits and sheds light on where interventions designed to reduce the harm of hate subreddits might most productively focus.
Limitations and Future Directions
Predictive modeling. The model we use is limited in its efficacy to predict where users will join. As mentioned previously, we cannot rule out the possibility that joining additional subreddits is a relatively stochastic decision or is affected by unobservable factors. Future work will need to explore motivations for joining additional hate subreddits in greater detail. Should a more predictive model become possible, however, this technology could help platforms develop proactive content moderation strategies by reducing problematic content on social media before it has been created (Habib et al. 2021). Moreover, while we employ techniques such as topic modeling to understand the textual features that are associated with users’ contributions to different hate groups, these techniques do not provide direct insight into why our model produced the predictions that it did. Existing methods and further advancements in the field of explainable natural language processing can help us understand how our predictive model learns to classify users as peripatetic or non-peripatetic (Zini and Awad 2022).
How peripatetic users contribute to subreddit bans. Previous research shows that the proportion of members of a subreddit who previously posted in banned or quarantined subreddits is strongly predictive of the subreddit being banned or quarantined in the future (Habib et al. 2021). In order to uncover the mechanisms by which peripatetic users contribute to given subreddits being banned, future work could more deeply investigate the activity of peripatetic users temporally in relation to the ban dates of each subreddit. Are these users simply interested in subreddits that are already likely to be banned, or does their activity within the subreddit directly or indirectly contribute to subreddit bans? Within each subreddit, one could identify large influxes of users who have already posted in other hate subreddits. Network analyses could be used to investigate the diffusion of ideas from peripatetic to non-peripatetic users, and dynamic topic modeling could be used to understand the shift in topics discussed by the core user bases of the subreddits over time (Wang, Blei, and Heckerman 2012). Such work could ultimately help to determine whether peripatetic users introduce problematic narratives to subreddits that eventually result in communities being banned.